Carrots cost $0.75 per kilogram and potatoes cost $0.70 per kilogram. A shopper bought 9 kilograms of these two vegetables for a toal of $6.60. How many kilograms of each did the shopper buy?
Answer by mananth(16946)

Carrots cost $0.75 per kilogram x kgs
potatoes cost $0.70 per kilogram. y kgs
9 kilograms of these two vegetables
for a total of $6.60.
x+y=9..........................1
0.75x+0.70y= 6.60
multiply by 100
75x+70y=660...................2
multiply equation 1 by -75 and add to eq. 2
-75x-75y=-675
75x+70y=660
-5y=-15
y=3 potatoes
x=6 carrots